OpenStudy (annjewell1998):

Please help me solve, i'm not looking for the answer, just the steps so i can get the answer. The formula for the amount with simple interest is A = P(1 + rt). Find t if A = 168, P = 150, and r = 0.08

OpenStudy (anonymous):

\[\frac{ A }{ P } = (1+rt)\] \[\frac{ A }{ P } -1 = rt\] \[\frac{ \frac{A}{P} -1 }{ r } = t\]

OpenStudy (annjewell1998):

okay so 168/150-1= 0.08t Right? I'm solving for t?

OpenStudy (annjewell1998):

Oh, 168/150-1/0.08= 1.5 therefore 1.5 is t? so 168/150-1=0.08(1.5) 12=12

OpenStudy (anonymous):

ANOTHER APPROACH: Plug in the given to the formula. \(\color{maroon}{A = P(1 + rt)}\) \(\color{olive}{A = 168}\), \(\color{indigo}{P = 150}\), and \(\color{orange}{r = 0.08}\) \(\huge \color{blue}{168=150(1+0.08t)}\) DISTRIBUTE 150 TO (1+0.08t). \(\huge \color{blue}{168=150+12t}\) SUBTRACT 150 ON BOTH SIDES. \(\huge \color{blue}{168-150=150-150+12t}\) \(\huge \color{blue}{18=12t}\) DIVIDE 12 ON BOTH SIDES. \(\huge \color{blue}{18 \div 12=12t \div 12}\) ------------------ \(\huge \color{maroon}{18 \div 12=t}\)

OpenStudy (annjewell1998):

Thanks to you both! :) this helped a lot.

OpenStudy (anonymous):

:)